At a particular temperature, N2O5 decomposes according to a first-order rate law with a half-life of 3.0 s. If the initial concentration of N2O5 is 1.0 × 1016 molecules/cm3, what will be the concentration in molecules/cm3 after 11.5 s?

A.

7.0 × 1014

B.

3.4 × 101

C.

1.0 × 1016

D.

2.0 × 1014

E.
Answer #1

Answer

for first order reaction  

rate constant (K)= 0.693/half life

rate constant (K)= 0.693/3 = 0.231 s^-1

now

for first order reaction

A= A0*e^-kt

here A= final concentration = ?

A0= initial concentration =1*10^16 molecules/cm^3

k= rate constant = 0.231 s^-1

t= time = 11.5 seconds

A= 1*10^16 *e^-0.231*11.5

A= 7.0*10^14 molecules/cm^3 (option A)